# WolBanking77: Wolof Banking Speech Intent Classification Dataset

### v1

@@ Abstract
Intent classification models have made a lot of progress in recent years. However, previous studies primarily focus on high-resource languages datasets, which results in a gap for low-resource languages and for regions with a high rate of illiterate people where languages are more spoken than read or written. This is the case in Senegal, for example, where Wolof is spoken by around 90% of the population, with an illiteracy rate of 42% for the country. Wolof is actually spoken by more than 10 million people in West African region. To tackle such limitations, we release a Wolof Intent Classification Dataset (WolBanking77), for academic research in intent classification. WolBanking77 currently contains 9,791 text sentences in the banking domain and more than 4 hours of spoken sentences. Experiments on various baselines are conducted in this work, including text and voice state-of-the-art models. The results are very promising on this current dataset. This paper also provides detailed analyses of the contents of the data. We report baseline f1-score and word error rate metrics respectively on NLP and ASR models trained on WolBanking77 dataset and also comparisons between models. We plan to share and conduct dataset maintenance, updates and to release open-source code.
@@ 1 Introduction
Wolof is spoken in Senegal, Gambia and Mauritania by more than 10 million people. Most of the population in Senegal speaks Wolof (around 90%)  [^1] https://www.axl.cefan.ulaval.ca/afrique/senegal.htm which is essentially a language of communication between several ethnic groups. More details on Wolof language are presented in the appendix section D.
@@
The lack of digital resources on Wolof motivated us to build the WolBanking dataset. Wolof, being an oral language, it is essential to establish voice assistants that allow the population access to digital services and to meet several challenges such as improving financial inclusion and access to digital public services. The trade sector, for instance, which plays an important role in the economy of African countries, is predominantly occupied by the informal sector [Ouattara et al., 2025]. We can also cite the agriculture, livestock, fishing, and transport sectors, which are all related to commercial activity and provide more jobs compared to the formal sector [Martínez and Short, 2022]. Recent advances in the field of artificial intelligence provide an excellent opportunity for these populations to benefit from the technology. They can gain better control over financial transactions and minimize the risk of fraud due to language barriers. In fact, the language barrier often forces business owners in the informal sector to trust a third party to consult their balance or make payments on their behalf. This potentially exposes them to the risk of fraud [Anthony et al., 2024, Ouattara et al., 2025].
@@
However, to build such a virtual assistant, it is necessary to understand the human request to be able to provide an appropriate response. In Natural Language Processing (NLP), the field that deals with the understanding of human language, is Natural Language Understanding (NLU). More particularly in our case, we propose to work with both text and speech given that our objective is to work with African languages which are based on oral tradition. Before doing an NLU task, the first step is to understand the information contained in speech using Spoken Language Understanding (SLU). According to the World Bank Bank [2022], the rate of adult illiterate population in Senegal is 42%, which means that to facilitate access to a virtual assistant for these populations, it is essential to offer voice processing solutions. To determine the intent of a user request, Coucke et al. [2018] approach the problem as an intent detection classification task, carried out from a text or an audio transcription.
@@
Models based on neural networks has emerged in recent years in the field of intent detection [Gerz et al., 2021, Krishnan et al., 2021, Si et al., 2023b]. However, very few datasets in African languages Conneau et al. [2022a], Mastel et al. [2023], Moghe et al. [2023], Mwongela et al. [2023], Kasule et al. [2024] have been explored. Even less in the field of e-banking for Sub-Saharan languages. Most datasets are in English Coucke et al. [2018] because development costs are very high, and substantial effort is required for their creation. This paper presents an intent classification dataset in Wolof, an African and Sub-Saharan language, based on the Banking77 dataset Casanueva et al. [2020a] which originally comprises both 13,083 customer service queries labeled with 77 intents and an audio dataset based on MINDS-14 Gerz et al. [2021] with 14 intents derived from a commercial e-banking system.
@@ 2.1 Existing datasets
Several datasets, both monolingual and multilingual, for NLU conversational question-answering system, have been published in the last decade. For instance, The Chatbot Corpus dataset composed of questions and answers and also The StackExchange Corpus based on the StackExchange platform both available on GitHub  [^2] https://github.com/sebischair/NLU-Evaluation-Corpora and evaluated by [Braun et al., 2017]. More recently, MINDS-14 Gerz et al. [2021] a multilingual dataset in the field of e-banking has been made publically available. When considering multilingual intent classification benchmarking, we can mention for instance the MultiATIS++ dataset Xu et al. [2020], which pertains to the aviation field, and has been expanded upon the original English ATIS dataset Price [1990] to six additional languages. More recently, the XTREME-S benchmark Conneau et al. [2022a] aims at evaluating several tasks such as speech recognition, translation, classification and retrieval. XTREME-S brings together several datasets covering 102 various languages, including 20 languages of Sub-Saharan Africa.
@@
Concerning the Wolof language, datasets have been published in the literature, particularly in the field of NLP and Automatic Speech Recognition (ASR). We can cite some datasets that contain Wolof texts such as afriqa dataset Ogundepo et al. [2023] which deals with the question answering (QA) task, masakhaner versions 1 and 2 Adelani et al. [2021, 2022b] which targets the Name Entity Recognition (NER) task, UD_Wolof-WTB  [^3] https://github.com/UniversalDependencies/UD_Wolof-WTB or even MasakhaPOS  [^4] https://github.com/masakhane-io/lacuna_pos_ner which deals with the part-of-speech tagging (POS) task. In addition, several datasets for the machine translation (MT) task such as OPUS,  [^5] https://opus.nlpl.eu/ FLORES 200 team et al. [2022], NTREX-128 Federmann et al. [2022] and MAFAND-MT [Adelani et al., 2022a]. Datasets in the field of ASR containing Wolof such as ALFFA Gauthier et al. [2016], fleurs Conneau et al. [2022b] and more recently KALLAAMA Gauthier et al. [2024] have been published. The cited datasets cover various fairly general areas from news or religion. However, to date, there is only one text dataset dedicated to the banking sector named INJONGO Yu et al. [2025] (with other domains such as home, kitchen and dining, travel and utility). This dataset includes slot-filling and intent classification for 16 African languages including Wolof.
@@
The authors Casanueva et al. [2020b] explored few-shot learning scenario in addition to the creation of the Banking77 dataset with 77 intents and 13,083 examples. ArBanking77 Jarrar et al. [2023] is the Arabic language version of the Banking77 dataset Casanueva et al. [2020a] in which the authors carried out simulations in low-resource settings scenario by training their model on a subset of the dataset. Other recent contributions to low-resource languages have been made, the authors Mastel et al. [2023] used Google Cloud Translation API to translate the ATIS dataset Price [1990] in Kinyarwanda and Swahili which are languages spoken by approximately 100 million people in East Africa. Moghe et al. [2023] publish a dataset called MULTI3NLU++ for several languages including Amharic. Kasule et al. [2024] propose a voice command dataset containing 20 intents in Luganda with the aim of being deployed on an IoT device.
@@ 2.2 Summary of contributions
In this work, 9,791 customer service queries from the Banking77 dataset Casanueva et al. [2020a] have been translated to French and Wolof by a team of linguistic experts from the Centre de Linguistique Appliquée de Dakar (CLAD). Additionally, this work releases another dataset based on MINDS-14 Gerz et al. [2021], with each query accompanied by audio data recorded from several speakers with varying accents and ages. The dataset covers 10 associated intentions with text and audio examples. The open source tool Lig-Aikuma Blachon et al. [2016] has been used to make the voice recordings. Our contributions include:
  • An audio dataset containing 263 sentences covering 10 intents in the banking and transport domains with varied voices and accents making this dataset the first of its kind in the Wolof language at the time of writing.
@@
  • Training and evaluation codes to facilitate the reproducibility of experimental benchmarks.
  • The results obtained with state-of-the-art models in various tasks such as Automatic Speech Recognition (ASR) and Intent Detection.
  • A dataset documention (datasheets) for WolBanking77.
@@ 3 Data collection
We present in this section the main information on the creation of WolBanking77, a more detailed description can be found in Appendix E.
@@ 3.1 Textual data
Duplicated sentences translated in Wolof has been removed for the Intent Detection task to avoid many-to-one translations from English to Wolof. Two versions of the dataset are reported on table 2. The first version is comprised of 5k samples which is a sub-sample of the second version of 9,791 samples. Train set is 80% of both versions and test set represents 20%. Note that intents are unbalanced, the most represented intent has a frequency of 200 while the least represented one has a frequency of 24.
@@ 3.2 Audio recordings and transcriptions
The audio corpus based on MINDS-14 Gerz et al. [2021] is made up of 186 queries and 77 responses initially translated from French into English thanks to a team of linguistic experts from the Centre de Linguistique Appliquée de Dakar (CLAD). The dataset also includes a phonetic transcription of the text in Wolof in order to facilitate the correct pronunciation of the sentences by the different speakers. Additional intents have been added to the initial MINDS-14 dataset which are OPEN_ACCOUNT for information when opening a bank account, BUS_RESERVATION for reserving a transport bus, TECHNICAL_VISIT for reserving an appointment for a technical inspection of a vehicle. TRANSFER_MONEY for the intention to transfer money and AMOUNT to specify the amount to transfer. See table 3.2 for the complete list of intents of the audio dataset.
@@
Figure 2 shows the top 5 most frequent words in the datastet after excluding stopwords. The dataset has 272 unique words and 3204 audio clips. The audio clips are in wav format using a single channel with a sampling rate of 16 kHz. Sentences have an average duration of 4815 ms. The intents ’CASH_DEPOSIT’, ’FREEZE’, ’LATEST_TRANSACTIONS’ contain the longest queries, while ’AMOUNT’, ’BUS_RESERVATION’, ’OPEN_ACCOUNT’ and ’TRANSFER_MONEY’ generally have a low query duration. The total duration of the dataset is around 4 hours and 17 minutes. Table 3: List of intents. intent domain BALANCE e-banking CASH_DEPOSIT FREEZE LATEST_TRANSACTIONS PAY_BILL OPEN_ACCOUNT TRANSFER_MONEY AMOUNT BUS_RESERVATION transport TECHNICAL_VISIT kontxaalisbëggoonajëflantexam65653737343428282727 Figure 2: Top 5 most frequent words: kont (account), xaalis (money), bëggoona (I wanted to), jëflante (operation), xam (know).
@@
The dataset was collected thanks to the participation of students from Cheikh Anta Diop University in Dakar (UCAD), more precisely at the Faculty of Letters and Human Sciences. Each participant recorded their voice using the elicitation mode of the Lig-Aikuma software [Blachon et al., 2016]. 186 utterances were pronounced by the participants in a closed environment. The elicitation mode of the Lig-Aikuma software was installed on an Android tablet. To start a session, information about the participant are requested such as native language, region of origin (see figure 3), gender, year of birth and name. This information is subsequently stored as metadata on the tablet’s memory card. More details on demographic data and distribution information can be found in Appendix A.1.
@@
To anonymize the name of the participants, the system generates a user_id which is used to replace the user name (more details on the ethical collection and processing of personal data are provided in Appendices A.6 and C). Then, a txt format file that contains the sentences to be pronounced is selected. One sentence at a time is presented during recording with the option to cancel recording if there is a mispronunciation at any stage. In terms of gender diversity, a total of 31 recording sessions were conducted including 14 males, 14 females and 3 unspecified. Text and audio data were split with 80% reserved for train set and 20% for test set. Before data splitting, some data preprocessing was made such as, removal of punctuation marks and numbers, lowercase text transformation. Corrupted audio files were removed from the dataset.
@@ 4.1 Automatic Speech Recognition
Automatic Speech Recognition (ASR) refers to the technology that enables a model to recognize and convert spoken language into text. ASR systems are widely used in various applications such as voice assistants, transcription services, customer service automation and more, whose researchers have done significant work such as Listen Attend and Spell by Chan et al. [2015] which is an end-to-end model for speech recognition based on a sequence-to-sequence neural network model. Graves et al. [2006] proposed Connectionist Temporal Classification (CTC) model used for training deep neural networks in speech recognition as well as other sequential problems where there is no explicit alignment information between the input and output. More recently, the NVIDIA team Żelasko et al. [2025] developed the Canary Flash model, which is one of the Canary Puvvada et al. [2024] variants and has the particularity of being multilingual and multitask. This model was able to achieve state-of-the-art results on several benchmarks including ASR.
Another state-of-the-art model published by Microsoft and named Phi-4-multimodal-instruct Microsoft et al. [2025] has recently been made available to address ASR tasks as well as in vision and text. Phi-4-multimodal shows an average word error rate score of 6.14% and occupies second place on Huggingface’s OpenASR leaderboard  [^6] https://huggingface.co/spaces/hf-audio/open_asr_leaderboard at the time of writing. To evaluate ASR models, Word Error Rate (WER) scores are reported in Section 5.4.

### v2

@@ Abstract
Intent classification models have made a significant progress in recent years. However, previous studies primarily focus on high-resource language datasets, which results in a gap for low-resource languages and for regions with high rates of illiteracy, where languages are more spoken than read or written. This is the case in Senegal, for example, where Wolof is spoken by around 90% of the population, while the national illiteracy rate remains at of 42%. Wolof is actually spoken by more than 10 million people in West African region. To address these limitations, we introduce the Wolof Banking Speech Intent Classification Dataset (WolBanking77), for academic research in intent classification. WolBanking77 currently contains 9,791 text sentences in the banking domain and more than 4 hours of spoken sentences. Experiments on various baselines are conducted in this work, including text and voice state-of-the-art models. The results are very promising on this current dataset. In addition, this paper presents an in-depth examination of the dataset’s contents. We report baseline F1-scores and word error rates metrics respectively on NLP and ASR models trained on WolBanking77 dataset and also comparisons between models. Dataset and code available at: wolbanking77.
@@ 1 Introduction
Wolof is spoken in Senegal, Gambia and Mauritania by more than 10 million people. Most of the population in Senegal speaks Wolof (around 90%)  [^1] https://www.axl.cefan.ulaval.ca/afrique/senegal.htm which is essentially a language of communication between several ethnic groups. More details on Wolof are presented in the appendix section 11.
@@
The lack of digital resources for Wolof motivated us to build the WolBanking77 dataset. Wolof, being an oral language, it is essential to establish voice assistants that allow the population access to digital services and help address challenges such as enhancing financial inclusion and access to digital public services. The trade sector, for instance, which plays an important role in the economy of African countries, is predominantly driven by the informal economy [Ouattara et al., 2025]. We can also cite the agriculture, livestock, fishing, and transport sectors, which are all related to commercial activity and generate more employment compared to the formal sector [Martínez and Short, 2022]. Recent advances in artificial intelligence provide an good opportunity for these populations to benefit from digital technology. They can gain better control over financial transactions and minimize the risk of fraud due to language barriers. In fact, the language barrier often compel business owners in the informal sector to rely on a third party to consult their account balances or make payments on their behalf. This potentially exposes them to the risk of fraud [Anthony et al., 2024, Ouattara et al., 2025].
@@
However, to build such a virtual assistant, it is necessary to understand human requests in order to provide appropriate responses. In Natural Language Processing (NLP), the field that deals with the understanding of human language, is Natural Language Understanding (NLU). In our case, we focus on both text and speech given that our objective is to work with African languages rooted in oral traditions. Before doing an NLU task, the first step is to understand the information contained in speech using Spoken Language Understanding (SLU). According to the World Bank Bank [2022], the rate of adult illiterate (see Appendix 8.4) population in Senegal is 42%, which means that to facilitate access to a virtual assistant for these populations, it is essential to offer voice processing solutions. To determine the intent of a user request, Coucke et al. [2018] frame the problem as an Intent Detection (ID) classification task, performed either directly from text or from audio transcriptions.
@@
Models based on neural networks has emerged in recent years in the field of ID [Gerz et al., 2021, Krishnan et al., 2021, Si et al., 2023b]. However, very few datasets in African languages Alexis et al. [2022], Mastel et al. [2023], Moghe et al. [2023], Mwongela et al. [2023], Kasule et al. [2024] have been explored. Even less in the field of e-banking for Sub-Saharan languages. Most existing datasets are in English Coucke et al. [2018], as their development requires significant financial and human resources. In this paper, we present an intent classification dataset in Wolof, an African and Sub-Saharan language, based on the Banking77 dataset Casanueva et al. [2020] which originally consists of 13,083 customer service queries labeled with 77 intents. In addition, our work also draws on the MINDS-14 dataset Gerz et al. [2021], which contains 14 intents derived from a commercial e-banking system and includes an audio component.
@@ 2.1 Existing datasets
Several datasets, both monolingual and multilingual, for NLU conversational question-answering system, have been published in the last decade. For instance, The Chatbot Corpus dataset composed of questions and answers and also The StackExchange Corpus based on the StackExchange platform both available on GitHub  [^2] https://github.com/sebischair/NLU-Evaluation-Corpora and evaluated by [Braun et al., 2017]. More recently, MINDS-14 Gerz et al. [2021] a multilingual dataset in the field of e-banking has been made publicly available. When considering multilingual intent classification benchmarking, we can mention for instance the MultiATIS++ dataset Xu et al. [2020], which pertains to the aviation field, and has been expanded upon the original English ATIS dataset Price [1990] to six additional languages. More recently, the XTREME-S benchmark Alexis et al. [2022] aims at evaluating several tasks such as speech recognition, translation, classification and retrieval. XTREME-S brings together several datasets covering 102 various languages, including 20 languages of Sub-Saharan Africa.
@@
With regard to the Wolof language, datasets have been published in the literature, particularly in the field of NLP and Automatic Speech Recognition (ASR). For example, some datasets include Wolof texts such as afriqa dataset Ogundepo et al. [2023] which deals with the question answering (QA) task, masakhaner versions 1 and 2 Adelani et al. [2021, 2022b] which targets the Name Entity Recognition (NER) task, UD_Wolof-WTB  [^3] https://github.com/UniversalDependencies/UD_Wolof-WTB or even MasakhaPOS  [^4] https://github.com/masakhane-io/lacuna_pos_ner which addresses the part-of-speech tagging (POS) task. In addition, several datasets for the machine translation (MT) task such as OPUS,  [^5] https://opus.nlpl.eu/ FLORES 200 team et al. [2022], NTREX-128 Federmann et al. [2022] and MAFAND-MT [Adelani et al., 2022a]. In the field of ASR, several datasets containing Wolof have also been published, including ALFFA Gauthier et al. [2016], fleurs Conneau et al. [2023] and more recently KALLAAMA Gauthier et al. [2024]. The cited datasets cover fairly general domains, such as news and religion. However, to date, there is only one text dataset dedicated to the banking sector named INJONGO Yu et al. [2025] (with other domains such as home, kitchen and dining, travel and utility). This dataset includes slot-filling and intent classification tasks for 16 African languages including Wolof.
@@
The authors Casanueva et al. [2020] explored few-shot learning scenario in addition to introducing the Banking77 dataset, which contains 77 intents and 13,083 examples. ArBanking77 Jarrar et al. [2023] is the Arabic language version of the Banking77 dataset Casanueva et al. [2020] in which the authors conducted simulations in low-resource settings scenario by training their model on a subset of the dataset. Other recent contributions to low-resource languages have been made, the authors Mastel et al. [2023] used Google Cloud Translation API to translate the ATIS dataset Price [1990] in Kinyarwanda and Swahili which are languages spoken by approximately 100 million people in East Africa. Similarly, Moghe et al. [2023] introduced the MULTI3NLU++ dataset for several languages including Amharic. Kasule et al. [2024] proposed a voice command dataset containing 20 intents in Luganda, designed for deployment on IoT devices.
@@ 2.2 Summary of contributions
In this work, 9,791 customer service queries from the Banking77 dataset Casanueva et al. [2020] was translated to French and Wolof by a team of linguistic experts from the Centre de Linguistique Appliquée de Dakar (CLAD). Additionally, this work introduces another dataset based on MINDS-14 Gerz et al. [2021], in which each query is paired with audio recordings from multiple speakers with diverse accents and ages (see Appendix 8.3). The dataset includes 10 intents represented with both text and audio examples. The open source tool Lig-Aikuma Blachon et al. [2016] was used to produce the voice recordings. Our contributions include:
  • The audio dataset contains 263 sentences covering 10 intents in the banking and transport domains, with diverse voices and accents, making it the first of its kind in the Wolof language at the time of writing.
@@
  • The results obtained with state-of-the-art models in various tasks such as Automatic Speech Recognition (ASR) and Intent Detection (ID). We also provide training and evaluation code to support the reproducibility of experimental benchmarks.
  • A dataset documentation (datasheets) for WolBanking77.
@@ 3 Data collection
In this section, we present the main information on the creation of WolBanking77, a more detailed description can be found in Appendix 12.
@@ 3.1 Textual data
Duplicated sentences translated in Wolof was removed for the ID task to avoid many-to-one translations from English to Wolof. Two versions of the dataset are reported on table 2. The first version is comprised of 5k samples which is a sub-sample of the second version of 9,791 samples. Train set is 80% of both versions and test set represents 20%. Note that intents are unbalanced, the most represented intent has a frequency of 200 while the least represented one has a frequency of 24.
@@ 3.2 Audio recordings and transcriptions
The audio corpus based on MINDS-14 Gerz et al. [2021] consists of 186 queries and 77 responses, initially translated from French into English by a team of linguistic experts from the Centre de Linguistique Appliquée de Dakar (CLAD). The dataset also includes a phonetic transcription of the Wolof text to facilitate the correct pronunciation of sentences by different speakers. Additional intents was added to the initial MINDS-14 dataset, namely: OPEN_ACCOUNT, for information related to opening a bank account, BUS_RESERVATION, for booking a transport bus (see Appendix 8.5), TECHNICAL_VISIT for scheduling a vehicle inspection appointment. TRANSFER_MONEY for the intent to transfer money and AMOUNT to specify the amount to be transferred. See table 3.2 for the complete list of intents in the audio dataset.
@@
Figure 2 shows top 5 most frequent words in the datastet after excluding stopwords. The dataset contains 272 unique words and 3,204 audio clips. The audio clips are in WAV format, single channel, with a sampling rate of 16 kHz. Sentences have an average duration of 4,815 ms. The intents ’CASH_DEPOSIT’, ’FREEZE’, ’LATEST_TRANSACTIONS’ contain the longest queries, whereas ’AMOUNT’, ’BUS_RESERVATION’, ’OPEN_ACCOUNT’ and ’TRANSFER_MONEY’ generally have shorter query durations. The total duration of the dataset is approximately 4 hours and 17 minutes. Table 3: List of intents. intent domain BALANCE e-banking CASH_DEPOSIT FREEZE LATEST_TRANSACTIONS PAY_BILL OPEN_ACCOUNT TRANSFER_MONEY AMOUNT BUS_RESERVATION transport TECHNICAL_VISIT kontxaalisbëggoonajëflantexam65653737343428282727 Figure 2: Top 5 most frequent words: kont (account), xaalis (money), bëggoona (I wanted to), jëflante (operation), xam (know).
@@
The dataset was collected with the participation of students from Cheikh Anta Diop University in Dakar (UCAD), specifically from the Faculty of Letters and Human Sciences. Each participant recorded their voice using the elicitation mode of the Lig-Aikuma software [Blachon et al., 2016]. A total of 186 utterances were recorded by participants in a controlled environment. The elicitation mode of the Lig-Aikuma software was installed on an Android tablet. At the start of each session, information about the participant are requested, including native language, region of origin (see figure 3), gender, year of birth and name. This information is subsequently stored as metadata on the tablet’s memory card. Note that, the scores presented in figure 3 are not representative of the population of each region. We tried to cover various ethnic groups in the country in order to have diverse accents and dialects as described in Appendix 8.2. Additional details on demographic data and distribution can be found in Appendix 8.1.
@@
To anonymize participant names, the system generates a user_id to replace the actual names (further details on the ethical collection and processing of personal data are provided in Appendices 8.11 and 10). Next, a text file containing the sentences to be pronounced is selected. During recording, sentences were presented one at a time, with the option to cancel if a mispronunciation occurred at any stage. In terms of gender diversity, a total of 31 recording sessions were conducted, including 14 males, 14 females and 3 unspecified. The text and audio data were split with 80% allocated for train set and 20% for test set. Prior to data splitting, preprocessing steps were applied, including the removal of punctuation marks and numbers, and conversion of text to lowercase. Corrupted audio files were also removed from the dataset.
@@ 4.1 Automatic Speech Recognition
Automatic Speech Recognition (ASR) refers to the technology that enables a model to recognize and convert spoken language into text. ASR systems are widely used in various applications such as voice assistants, transcription services and customer service automation. Significant research has been conducted in this area, including Listen Attend and Spell by Chan et al. [2016], an end-to-end speech recognition system based on a sequence-to-sequence neural network. Graves et al. [2006] proposed Connectionist Temporal Classification (CTC) model used for training deep neural networks in speech recognition as well as other sequential problems where there is no explicit alignment information between the input and output. More recently, the NVIDIA team Żelasko et al. [2025] developed the Canary Flash model, a variant of Canary models Puvvada et al. [2024] notable for being both multilingual and multitask. This model achieved state-of-the-art results on several benchmarks, including ASR.
Another state-of-the-art model developed by Microsoft and named Phi-4-multimodal-instruct Microsoft et al. [2025] has recently been released to address ASR tasks, as well as vision and text applications. Phi-4-multimodal achieves an average WER score of 6.14% and currently ranks second on Huggingface’s OpenASR leaderboard  [^6] https://huggingface.co/spaces/hf-audio/open_asr_leaderboard. To evaluate ASR models, Word Error Rate (WER) scores are reported in Section 5.4.